# LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

### v1

@@ Appendix C Models Inference and Training Details
(1) Predicted label: group-based solidarity towards expellees
@@
?[…] das Lastenausgleichsgesetz und die Eingliederung der Vertriebenen [steht] in keinem Widerspruch zu unserm Anspruch und unserm Recht auf die Heimat. Im Gegenteil, […] daß die Rücksiedlung […] nur durchgeführt werden kann, wenn die körperliche, seelische und auch die wirtschaftliche Substanz der Vertriebenen erhalten geblieben ist, und das kann nur geschehen, wenn sie in die westdeutsche Wirtschaft eingegliedert werden. […] Wir sind durchaus für die Ausgestaltung des Lastenausgleichs nach sozialen Gesichtspunkten; […] aber der Rechtsanspruch darf nicht so behandelt werden, als sei er nur etwas Zweitrangiges, angewiesen auf das, was übrigbleibt.? ?[…] the Equalization of Burdens Act and the integration of the expellees does not contradict our right to the homeland. On the contrary, resettlement […] can only be achieved if the physical, mental, and economic substance of the expellees has been preserved, and that can only happen if they are integrated into the West German economy. […] We are fully in favor of shaping the Equalization of Burdens in accordance with social considerations; […] but the legal entitlement should not be treated as secondary, dependent only on what is left over.?
(2) Predicted label: exchange-based solidarity towards expellees
@@
?Ich darf […] auf einige Gesichtspunkte hinweisen […] zum Bundesvertriebenengesetz. Der Ihnen vorliegende Entwurf geht […] von der Zielsetzung aus, die […] Leistungskraft der Vertriebenen und Flüchtlinge […] im Rahmen der Gesamtwirtschaft zur Geltung zu bringen. Die Eingliederungsbestimmungen sollen der Erfüllung berechtigter Ausgleichsansprüche hinsichtlich der verlorenen Existenz dienen […] mit dem Ziel, daß sich das gesamte Arbeitspotential, das der Einheimischen und der Vertriebenen, so betätigen kann, daß eine größtmögliche Wirkung erzielt wird.? ?I would like to point out some aspects […] on the Federal Expellees Act. The draft before you starts from the objective […] of making […] productive capacities of the expellees and refugees […] effective within the overall economy. The integration provisions are intended to meet legitimate compensation claims with regard to lost livelihoods […] with the goal that the entire labor potential, that of natives and of expellees alike, can be deployed in such a way that the greatest possible effect is achieved.?
(3) Predicted label: compassionate solidarity towards Soviet-zone refugees
@@
(4) Predicted label: group-based anti-solidarity towards expellees
@@
Table 7: Examples for postwar period discussed in Section 4.2. Bold text is the main sentence, the other sentences are for context.
@@
(5) Predicted label: group-based solidarity towards Soviet-zone refugees
@@
?[…] den Ländern [wurden] für die wohnungsmäßige Unterbringung der […] Sowjetzonenflüchtlinge […] 350 Millionen DM […] zur Verfügung gestellt […]. Der Bundestag ist […] davon ausgegangen, daß die Sowjetzonenflüchtlinge im Lastenausgleich nicht berücksichtigt werden konnten und daß die Verweisung […] auf den Härtefonds nur als Interimslösung gedacht war. […] Die Bundesregierung prüft […] noch […] Maßnahmen, um den bei den Sowjetzonenflüchtlingen auftretenden Notständen abzuhelfen.? ?[…] the federal government provided 350 million DM for the housing of […] Soviet-zone refugees […]. Parliament assumed that Soviet-zone refugees could not be included in the Equalization of Burdens and that referral […] to the hardship fund was intended only as an interim solution. […] The federal government is still examining measures to address the emergencies arising among the Soviet-zone refugees.?
(6) Predicted label: group-based anti-solidarity towards foreigners
@@
?[…] Wir müssen das Bundesgebiet nicht nur gegen politisch subversive Elemente, wir müssen uns auch gegen kriminelle Elemente abschirmen. […] Für die Bundesrepublik kommt […] hinzu, daß sie […] Anlaß hat, sich gegen Zehntausende von heimatlosen Ausländern abzuschirmen, die […] in der Absicht [kommen], bei uns wirtschaftliche Versorgung oder politisches Asyl zu finden. […] alle internationalen Verpflichtungen, die wir […] erfüllen, [gehören] selbstverständlich nicht unter das, was ich hier ausgeführt habe.? ?[…] We must shield the Federal territory not only against politically subversive elements, we but also against criminal elements. […] For the Federal Republic there is the added circumstance that it has reason to shield itself against tens of thousands of stateless foreigners who […] come with the intention of seeking economic support or political asylum here. […] all international obligations which we […] fulfill, of course, do not fall under what I have set out here.?
(7) Predicted label: group-based solidarity towards foreigners
@@
?[…] Auch in der öffentlichen Fürsorge sollen sie Leistungen in gleicher Höhe erhalten wie deutsche Staatsangehörige. Es soll ihnen Zutritt zu Schulen aller Art zu den gleichen Bedingungen wie Deutschen gewährt werden. […] die Bundesrepublik [gewährt] ihnen einen Rechtsschutz, dessen sie bedürfen, weil sie nicht mehr den Schutz ihres Heimatstaates genießen […]. Der Gesetzentwurf geht […] über den Konventionsentwurf, dessen Grundsätze Mindestforderungen für die Behandlung heimatloser Ausländer darstellen, hinaus. […]? ?[…] In public welfare too, they are to receive benefits equal to those of German citizens. They are to be granted access to schools of all kinds under the same conditions as Germans. […] the Federal Republic shall grant them legal protection, which they need because they no longer enjoy the protection of their home state […]. The draft law goes […] beyond the draft convention, whose principles set minimum standards for the treatment of stateless foreigners. […]?
(8) Predicted label: group-based solidarity towards foreigners
@@
(9) Predicted label: empathic solidarity towards foreigners
@@
Table 8: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(10) Predicted label: exchange-based solidarity towards expellees,
@@
(11) Predicted label: group-based solidarity towards expellees,
@@
(12) Predicted label: compassionate solidarity towards refugees,
@@
(13) Predicted label: group-based anti-solidarity towards expellees,
@@
Table 9: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(14) Predicted label: group-based solidarity towards refugees,
@@
Table 10: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
@@
(15) Predicted label: compassionate solidarity towards refugees

### v2

@@ Appendix E Label Distribution Distance Metrics
[…]
? […] das Interessante […] ist, daß nicht nur die östliche Landwirtschaft sondern auch die Industrie Ausländer braucht, daß die Zahlen der industrietätigen Ausländer im Lause der Jahre viel mehr gewachsen sind als die Zahlen der von der Landwirtschaft gebrauchten Arbeiter. […] Das Maß des Bedürfnisses der Beschäftigung von Ausländern ist natürlich auch hier örtlich und nach Betrieben verschieden. Es gibt gewisse Betriebe, für die inländische Arbeiter kaum noch zu bekommen sind. […] Es gibt aber auch industrielle Gebiete, die ohne diese Ausländer gar nicht weiter existieren können. ? ? […] what is interesting […] is that not only eastern agriculture but also industry needs foreigners, and that the number of foreigners employed in industry has grown over the years much more than the number of workers needed in agriculture. […] The extent of the need for the employment of foreigners naturally also varies here by locality and by sector. There are certain sectors in which domestic workers can scarcely be found anymore. […] There are also industrial regions that simply could not continue to exist without these foreigners. ?
[…]
?[…] Die Beschleunigung der Asylverfahren ist überfällig, und eine konsequente Rückführung […], und […] eine konsequente Abschiebung. Übrigens sind zügige Verfahren auch und gerade im Interesse der Menschen mit Bleibeperspektive, die wir dann umso besser integrieren können, weil wir ihnen eine Perspektive geben können. Zu guter Letzt müssen wir die Einwanderung nach Deutschland insgesamt neu ordnen, mit einem Einwanderungsgesetz, das diesen Namen verdient. Herr Kollege Castellucci, möglicherweise wird umgekehrt ein Schuh daraus: Wer Akzeptanz für legale Migration erhöhen will, der muss auch entschlossen illegale Migration bekämpfen. […]? ?Speeding up asylum procedures is overdue, as are consistent returns […] and […] consistent deportations. Swift procedures are also, and especially, in the interest of people who have a prospect of staying, because we can then integrate them all the better by giving them a perspective. Finally, we must reorganize immigration to Germany as a whole, with an immigration law worthy of the name. Mr Castellucci, perhaps the reverse is true: anyone who wants to increase acceptance of legal migration must also resolutely combat illegal migration. […]?
[…]
?[…] Sie [die europäischen Staaten] sind aber nicht bereit, Flüchtlinge vor Menschenrechtsverletzungen zu schützen, sondern diffamieren sie als Schmarotzer und Bedrohung unserer Sicherheit, wie dies leider auch gerade durch den Kollegen der SPD wieder geschehen ist. Der Vertrag von Maastricht sieht die gemeinsame Einführung eines Visumzwangs vor, wenn auf Grund einer Notlage in einem Land eine größere Fluchtbewegung zu erwarten ist. So wird ganz bewußt verfolgten oder unter sonstigen Notlagen leidenden Menschen die Flucht erheblich erschwert, oft sogar unmöglich gemacht. Flüchtlinge werden damit geradezu in die Arme von Schlepperorganisationen getrieben, über die man sich hier so gerne beschwert.? ?[…] But they [the European states] are not willing to protect refugees from human rights violations; instead, they defame them as parasites and a threat to our security, as unfortunately just happened again in the remarks of the SPD colleague. The Maastricht Treaty provides for the joint introduction of a visa requirement if, because of an emergency in a country, a larger refugee movement is to be expected. In this way, people who are persecuted or otherwise suffering acute hardship are deliberately making it much harder to flee, often even impossible. Refugees are thereby virtually driven into the arms of smuggling organizations, about which people here so readily complain.?
Table 11: Error analysis examples. Bold text is the main sentence, the other sentences are for context. The reference label is the final human consensus label. Where relevant, we also report the distribution of individual human annotations.
Original Text Translation
Example (4): Predictions: group-based solidarity (gpt-oss, Qwen), group-based anti-solidarity (Llama). Human reference: group-based solidarity (10 Apr 2024)
?Kommen wir zum Entwurf der AfD und zu dem, was ihn einfach grottenschlecht macht. Sie fordern, dass in allen Fällen, in denen ein Elternteil nichtdeutscher Herkunft ist, die Ausländerbehörde beteiligt wird. Nicht nur stellen Sie damit alle Ausländer/-innen unter Generalverdacht […], sondern Sie verstoßen damit auch gegen EU-Recht […]. Sie ?übersehen? in Ihrem Entwurf, dass ein Großteil der Ausländer, die hier die Vaterschaftsanerkennung beantragen, längst eine Niederlassungserlaubnis haben. Aber gut, dass Sie gegen unsere ausländischen Mitbürgerinnen und Mitbürger hetzen. Ist ja wirklich Ihr politisches Tagesgeschäft, so bedauerlich das auch ist. […] ? ?[…] Let us turn to the AfD’s draft and to what makes it simply atrocious. You demand that in all cases where one parent is of non-German origin, the immigration authority must be involved. Not only does this place all foreigners under general suspicion […], but it also violates EU law […]. You overlook in your draft that a large proportion of the foreigners who apply here for recognition of paternity already have permanent residency. But of course, it fits your agenda to agitate against our foreign fellow citizens — that is really your daily political business, as regrettable as it is. […]?
[…]
? […] Diese Rückführung [der Evakuierten] muß beschleunigt werden, und wir müssen gesetzgeberische Maßnahmen ergreifen, damit die Evakuierten nicht immer wieder auf lange Zeit vertröstet werden, sondern in absehbarer Zeit zu ihren Rechten, zu einer menschenwürdigen Behandlung kommen. […] Gerade das Vertrauen dieser Millionen von Menschen verpflichtet uns, sie auch in den nächsten Jahren nicht zu enttäuschen, […] diesen Eingliederungsprozeß auf allen Gebieten mit dem größten Nachdruck zu fördern […]. In der politischen Diskussion hat zwar das Problem der sachlichen, wirtschaftlichen und sozialen Interessen der Vertriebenen und aller Kriegsgeschädigten kaum noch eine Rolle gespielt […]. ? ? […] This return [of evacuees] must be accelerated, and we must take legislative measures so that evacuees are not kept waiting indefinitely time and again, but instead obtain their rights and humane treatment within a foreseeable period of time. […] It is precisely the trust of these millions of people that obliges us not to disappoint them in the years ahead either, […] and to promote this process of integration in every sphere with the greatest possible determination […]. In political debate, the problem of the material, economic, and social interests of the expellees and all other victims of war has in fact hardly played any role any longer […]. ?
[…]
?Die Bundesrepublik hat […] einen Vorbehalt zur UN-Kinderrechtskonvention zurückgenommen, der ihr zusichern sollte, dass sie ausländische Kinder schlechter behandeln kann als inländische. […] In den vorliegenden Anträgen […] ist dargelegt, was alles zu tun wäre: Heraufsetzung der Verfahrensmündigkeit von minderjährigen Flüchtlingen von 16 auf 18 Jahre, Ausbau und Verbesserung der Betreuung unbegleiteter minderjähriger Flüchtlinge durch die Jugendfürsorge, ihre Herausnahme aus dem Flughafenverfahren, keine Abschiebehaft für Minderjährige, keine Unterbringung von Kindern und Jugendlichen in Sammelunterkünften für Asylsuchende und Flüchtlinge. Unbegleitete Kinder und Jugendliche, deren Eltern nicht ausfindig gemacht werden können, sollen genauso behandelt werden wie elternlose deutsche Kinder. […] ? ?The Federal Republic has […] withdrawn a reservation to the UN Convention on the Rights of the Child that had ensured it could treat foreign children worse than German children. […] The motions submitted […] outline what still needs to be done: raising the age of legal capacity in proceedings for minor refugees from 16 to 18; expanding and improving youth welfare support for unaccompanied minor refugees; removing them from the airport procedure; no detention pending deportation for minors; no accommodation of children and adolescents in collective facilities for asylum seekers and refugees. Unaccompanied children and adolescents whose parents cannot be located should be treated in the same way as German children without parents. […] ?
Table 12: Error analysis examples. Bold text is the main sentence, the other sentences are for context. The reference label is the final human consensus label. Where relevant, we also report the distribution of individual human annotations.
Original Text Translation
[…]
?Das Hamburger Modell ist geradezu vorbildlich; denn hier werden nicht Menschen mit geringem Einkommen und Geflüchtete gegeneinandergestellt, sondern alle haben den gleichen Anspruch auf Mobilität, mit dem Unterschied, dass die Geflüchteten das von ihrem Taschengeld zwangsweise abgezogen bekommen. […] ?Wir sind auf Migration angewiesen, aber die Flüchtlinge…?, könnte man sagen. […] Ich bin so erzogen: Wenn jemand Hilfe braucht und du kannst helfen, dann hilf. Ich bin stolz und froh, dass unser Land helfen kann […].? ?The Hamburg model is virtually exemplary, because here people with low incomes and refugees are not being pitted against one another; rather, everyone has the same entitlement to mobility, with the difference that refugees have it forcibly deducted from their allowance. […] One might say, ?We depend on migration, but the refugees…?. […] I was raised to believe this: if someone needs help and you can help, then help. I am proud and glad that our country is able to help […].?
[…]
?[…] Die große Aufgabe, die uns hier gesetzt war, war der Versuch, einen Kompromiß […] zwischen den Interessen der Ausländer in der Bundesrepublik Deutschland und den berechtigten Ansprüchen der deutschen Bevölkerung auf Aufrechterhaltung der öffentlichen Ordnung und Sicherheit zu finden. Das Gesetz ist ein Ausdruck unserer freiheitlichen demokratischen Grundordnung, indem es den Ausländern in der Bundesrepublik ein sehr hohes Maß von Entfaltungsmöglichkeiten gewährt. […] Das Gesetz nimmt auch besonders Bedacht auf eine Verbesserung der Rechtslage der ausländischen Flüchtlinge […], die ihre Heimat unter dem Druck totalitärer Systeme verlassen mußten. […]? ?[…] The major task before us was to try to find a compromise […] between the interests of foreigners in the Federal Republic of Germany and the legitimate claims of the German population to the maintenance of public order and security. This law is an expression of our liberal democratic constitutional order in that it grants foreigners in the Federal Republic a very high degree of scope for self-development. […] The law also gives special consideration to improving the legal position of foreign refugees […] who were forced to leave their homeland under pressure from totalitarian systems. […]?
[…]
?[…] Mit diesem auf Angebotsorientierung im Migrationsrecht zielenden Gesetzentwurf wollen die Grünen im Kern nichts anderes, als die Einwanderung massiv vereinfachen und die Tore weit aufstoßen. Wir als Unionsfraktion stellen uns gegen die Beliebigkeit im Einwanderungsrecht und stehen daher auch klar für eine gezielte Auswahl bei der Migration ein. Der Staat und die Politik haben die Pflicht, bewusst darauf zu achten, dass die Migration für unser Land vor allem einen Mehrwert und keine Belastung darstellt, und genau darin liegt der Unterschied zum Asyl- und Flüchtlingsrecht. […] ? ?[…] With this bill, which is aimed at a supply-oriented approach to migration law, the Greens essentially want nothing other than to make immigration much easier and throw the gates wide open. As the CDU/CSU parliamentary group, we oppose arbitrariness in immigration law and therefore clearly support a targeted selection in migration. The state and policymakers have a duty to ensure consciously that migration represents above all an added value for our country and not a burden, and that is precisely the difference from asylum and refugee law. […]?
Table 13: Error analysis examples. Bold text is the main sentence, the other sentences are for context. The reference label is the final human consensus label. Where relevant, we also report the distribution of individual human annotations.
Original Text Translation
[…]
?52 Prozent der britischen Muslime sagen nicht nur: ?Homosexualität ist moralisch falsch?, sondern: Homosexualität sollte verboten werden. […] bei fortgesetzter Einwanderung [wird sich] der Anteil der Muslime in Deutschland in den kommenden Jahren auf 20 Prozent erhöhen. In vielen Städten und Gemeinden sind sie dann die Mehrheit. Der Islam kennt keine Minderheitenrechte. Das sind die dunklen Wolken, die über den Homosexuellen in Westeuropa aufziehen.? ?52 percent of British Muslims say not only, ?Homosexuality is morally wrong?, but also: homosexuality should be banned. […] if immigration continues, the share of Muslims in Germany will rise to 20 percent in the coming years. In many cities and municipalities, they will then be the majority. Islam recognizes no minority rights. These are the dark clouds gathering over homosexuals in Western Europe.?
[…]
?Ausländer bleibt unerachtet solchen Aufenthalts innerhalb der Mecklenburgischen Lande immer Ausländer, und einem solchen die Trauung innerhalb der diesseitigen Lande zu gestatten, ist allein das Großherzogliche Ministerium des Innern competent. […] Vielmehr muß in solchen Fällen der Wiederverheirathung der Mecklenburger sein Recht zur Verheiratung durch einen gesetzlichen Domizilschein über das ihm für einen Mecklenburgischen Ort ertheilte Recht der Niederlassung und Verheiratung nachweisen und der Ausländer den Konsens des Großherzoglichen Ministeriums des Innern beibringen, bevor die Trauung geschehen darf. Ja, meine Herren, da sehen Sie nun, was in Mecklenburg das Indigenat aus dem Artikel 2 der Bundesverfassung für eine Bedeutung hat, und wie nothwendig es ist, daß irgend eine Ergänzung des Freizügigkeitsgesetzes […] erfolge, damit das Gesetz dort nicht blos eine Schale ohne Kern bleibe.? ?A foreigner remains a foreigner within the lands of Mecklenburg despite such residence, and only the Grand Ducal Ministry of the Interior is competent to permit such a person to marry within these territories. […] Rather, in such cases of remarriage, the Mecklenburger must prove his right to marry by means of a lawful certificate of domicile showing that he has been granted the right of residence and marriage in a Mecklenburg locality, and the foreigner must produce the consent of the Grand Ducal Ministry of the Interior before the marriage may take place. Gentlemen, there you can see what significance indigeneity under Article 2 of the federal constitution has in Mecklenburg, and how necessary it is that some supplementation of the Freedom of Movement Act […] should be made, so that the law there may not remain merely a shell without a kernel.?
[…]
?Wann und warum ein Ausländer ohne ein vorangegangenes gerichtliches Erkenntniß aus dem Reich oder einem Partikularstaat ausgewiesen werden darf, darüber existiren zwar keine gesetzlichen Normen und Bestimmungen […]. Rechtfertigende Gründe der Ausweisung eines Ausländers nach dem allgemein anerkannten Völkerrechte sind folgende: erstens Gefährdung der öffentlichen Sicherheit durch ein feindliches Auftreten des Ausgewiesenen […]; zweitens […]; drittens[…], und viertens […]. Damit sind alle Fälle erschöpft, in welchen nach dem modernen Völkerrecht Ausweisungen von einzelnen Personen stattfinden dürfen, und alle Lehrer des Völkerrechts sind um so vorsichtiger in der Empfehlung auch nur eines dieser Mittel, als sich das moderne Völkerrecht immer mehr der Theorie zuwendet, daß ein domizilirter Ausländer zu den wirklichen Staatsangehörigen gehört und als solcher zu behandeln ist. […]? ?Although there are no legal norms or provisions governing when and why a foreigner may be expelled from the Empire or from an individual state without prior judicial judgment […], the reasons justifying the expulsion of a foreigner under generally recognized international law are the following: first, endangerment of public security through hostile conduct on the part of the person expelled […]; second […]; third […]; and fourth […]. These exhaust all the cases in which, under modern international law, expulsions of individual persons may take place, and all teachers of international law are all the more cautious in recommending even one of these measures because modern international law is increasingly turning toward the theory that a domiciled foreigner belongs among the true members of the state and is to be treated as such. […]?
Table 14: Error analysis examples. Bold text is the main sentence, the other sentences are for context. The reference label is the final human consensus label. Where relevant, we also report the distribution of individual human annotations.
Original Text Translation
@@
(1) Dominant DSL subtype score: group-based solidarity towards expellees
@@
?[…] das Lastenausgleichsgesetz und die Eingliederung der Vertriebenen [steht] in keinem Widerspruch zu unserm Anspruch und unserm Recht auf die Heimat. Im Gegenteil, […] daß die Rücksiedlung […] nur durchgeführt werden kann, wenn die körperliche, seelische und auch die wirtschaftliche Substanz der Vertriebenen erhalten geblieben ist, und das kann nur geschehen, wenn sie in die westdeutsche Wirtschaft eingegliedert werden. […] Wir sind durchaus für die Ausgestaltung des Lastenausgleichs nach sozialen Gesichtspunkten […]. ? ?[…] the Equalization of Burdens Act and the integration of the expellees does not contradict our right to the homeland. On the contrary, resettlement […] can only be achieved if the physical, mental, and economic substance of the expellees has been preserved, and that can only happen if they are integrated into the West German economy. […] We are fully in favor of shaping the Equalization of Burdens in accordance with social considerations […]. secondary, dependent only on what is left over. ?
(2) Dominant DSL subtype score: group-based solidarity towards expellees
@@
?[…] Der Ihnen vorliegende Entwurf [zum Bundesvertriebenengesetz] geht […] von der Zielsetzung aus, die […] Leistungskraft der Vertriebenen und Flüchtlinge […] im Rahmen der Gesamtwirtschaft zur Geltung zu bringen. Die Eingliederungsbestimmungen sollen der Erfüllung berechtigter Ausgleichsansprüche hinsichtlich der verlorenen Existenz dienen […] mit dem Ziel, daß sich das gesamte Arbeitspotential, das der Einheimischen und der Vertriebenen, so betätigen kann, daß eine größtmögliche Wirkung erzielt wird.? ?[…] The draft [on the Federal Expellees Act] before you starts from the objective […] of making […] productive capacities of the expellees and refugees […] effective within the overall economy. The integration provisions are intended to meet legitimate compensation claims with regard to lost livelihoods […] with the goal that the entire labor potential, that of natives and of expellees alike, can be deployed in such a way that the greatest possible effect is achieved.?
(3) Dominant DSL subtype score: compassionate solidarity towards Soviet-zone refugees
@@
(4) Dominant DSL subtype score: group-based anti-solidarity towards expellees
@@
Table 15: Examples for postwar period discussed in Section˜4.2. Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
@@
(5) Dominant DSL subtype score: group-based/compass. solidarity towards Soviet-zone refugees
@@
?[…] [Sowjetzonenflüchtlinge] meinen, es könnte ihnen sonst bei mangelnder Bereitwilligkeit der Bevölkerung zu neuen Opfern vielleicht nicht einmal mehr das zur Verfügung stehen, was ihnen heute aus dem Härtefonds zur Verfügung steht. […] Wir halten es für unvertretbar, daß […] diese nach wie vor auf den genannten, im Rahmen des Gesamtlastenausgleichs verhältnismäßig geringen Betrag ohne Rechtsanspruch angewiesen sein sollen. [Die], deren Urstatus dem der Heimatvertriebenen in Westdeutschland doch gleich ist, im Lastenausgleich diesen Heimatvertriebenen nicht gleichgestellt sind.? ?[…] [Soviet-zone refugees] fear that, if public willingness for further sacrifices declines, they may no longer receive even the support currently provided through the hardship fund […]. [It is] indefensible that […] they should still be dependent on the aforementioned, comparatively small amount within the overall equalization scheme without any legal entitlement. […] those whose original legal status is equivalent to that of expellees in West Germany are not treated equally under the Equalization of Burdens system.?
(6) Dominant DSL subtype score: group-based anti-solidarity towards foreigners
@@
?[…] Wir müssen das Bundesgebiet nicht nur gegen politisch subversive Elemente, wir müssen uns auch gegen kriminelle Elemente abschirmen. […] die Bundesrepublik […] [hat] Anlaß, sich gegen Zehntausende von heimatlosen Ausländern abzuschirmen, die […] in der Absicht [kommen], bei uns wirtschaftliche Versorgung oder politisches Asyl zu finden. […] alle internationalen Verpflichtungen, die wir […] erfüllen, [gehören] selbstverständlich nicht unter das, was ich hier ausgeführt habe.? ?[…] We must shield the Federal territory not only against politically subversive elements, we but also against criminal elements. […] the Federal Republic […] has reason to shield itself against […] stateless foreigners who […] come with the intention of seeking economic support or political asylum here. […] all international obligations which we […] fulfill, of course, do not fall under what I have set out here.?
(7) Dominant DSL subtype score: group-based solidarity towards foreigners
@@
?[…] [HAuslG] […] [wird] den Besatzungsmächten den Entschluß erleichtern, der Bundesrepublik die volle Verantwortung für die heimatlosen Ausländer zu übertragen. Wenn […] das deutsche Volk seine Bereitschaft bekundet, die Eingliederung der heimatlosen Ausländer in die deutsche Betreuung […] zu erleichtern, so gehen wir dabei von der Erwartung aus, daß auch die heimatlosen Ausländer bemüht sein werden, sich in die nun einmal gegebene Lage einzufügen. Sie werden sich dabei bewußt sein müssen, daß auch die Lage der deutschen Bevölkerung schwierig ist und daß viele Zehntausende insbesondere von Heimatvertriebenen noch immer arbeitslos sind und über keinen angemessenen Wohnraum verfügen. […] der heimatlose Ausländer [wird] viele der Härten dieser Nachkriegszeit mit den Deutschen im Gebiete der Bundesrepublik teilen müssen. ? ? […] [HAuslG] […] will make it easier for the occupying powers to decide to transfer full responsibility for stateless foreigners to the Federal Republic. If […] the German people declare their willingness to facilitate […] the integration of stateless foreigners into German administration and care, we proceed on the assumption that the stateless foreigners will also strive to adapt to the situation as it now stands. They must remain aware that the situation of the German population is also difficult and that many tens of thousands, especially expellees, remain unemployed and lack adequate housing. […] stateless foreigners will have to share many of the hardships of this postwar period with the Germans in the territory of the Federal Republic. ?
(8) Dominant DSL subtype score: group-based solidarity towards foreigners
(Oct 18, 1950)
?[…] Auch in der öffentlichen Fürsorge sollen sie Leistungen in gleicher Höhe erhalten wie deutsche Staatsangehörige. Es soll ihnen Zutritt zu Schulen aller Art zu den gleichen Bedingungen… gewährt werden. […] die Bundesrepublik [gewährt] ihnen einen Rechtsschutz, dessen sie bedürfen, weil sie nicht mehr den Schutz ihres Heimatstaates genießen […]. Der Gesetzentwurf geht […] über den Konventionsentwurf, dessen Grundsätze Mindestforderungen für die Behandlung heimatloser Ausländer darstellen, hinaus. […]? ?[…] In public welfare too, they are to receive benefits equal to those of German citizens. They are to be granted access to schools of all kinds under the same conditions […]. the Federal Republic shall grant them legal protection, which they need because they no longer enjoy the protection of their home state. The law goes […] beyond the draft convention, whose principles set minimum standards for the treatment of stateless foreigners. […]?
Table 16: Examples for postwar period discussed in Section˜4.2 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(9) Dominant DSL subtype score: exchange-based solidarity towards foreigners
@@
(10) Dominant DSL subtype score: group-based solidarity towards foreigners
@@
(11) Dominant DSL subtype score: exchange-based solidarity towards expellees,
@@
(12) Dominant DSL subtype score: group-based/
compassionate solidarity towards expellees,
@@
Table 17: Examples for postwar period discussed in Section˜4.2 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(13) Dominant DSL subtype score: compassionate solidarity towards refugees,
@@
(14) Dominant DSL subtype score: group-based anti-solidarity towards expellees,
@@
(15) Dominant DSL subtype score: group-based solidarity towards refugees,
@@
(16) Dominant DSL subtype score: exchange-based solidarity towards refugees
Party: CDU/CSU
(Jan 31, 1951)
?[…] Es ist Voraussetzung, dass das deutsche Volk für das System des Lastenausgleichs einen Weg findet, damit das, was aus der deutschen Produktion und Wirtschaftskraft herausgenommen werden muss, […] dafür verwendet wird, die Kräfte der Kriegsgeschädigten, der Flüchtlinge und der Heimatvertriebenen wieder in die deutsche Produktion hineinzustellen und diese damit zu stärken. […] Was wir an Eingliederungshilfe […], an Wohnraumhilfe […] und an Förderungshilfe aller Art […] verwenden, ist […] produktive Ausgabe. […] das ist rund die Hälfte des gesamten Aufkommens. ? ?[…] It is a prerequisite that the German people find a way, within the system of burden sharing, to ensure that what must be taken from German production and economic strength is used […] to restore the productive capacity of war victims, refugees, and expellees and to reintegrate them into German production, thereby strengthening it. […] What we allocate for integration assistance […], for housing assistance […], for all kinds of development assistance is […] productive expenditure. […] that is roughly half of the total revenue.?
Table 18: Examples for postwar period discussed in Section˜4.2 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
@@
(17) Dominant DSL subtype score: group-based anti-solidarity towards expellees,
Party: DP
(March 18, 1953)
?Also auch der dritte, vierte und fünfte Sohn eines ostdeutschen Bauern […] kann alle […] Vergünstigungen erhalten. […] Welcher einheimische Bauer kann […] seinen drei oder vier Jungen eine solche Mitgift erarbeiten? Wir schaffen, wenn wir die alte Vorlage wiederherstellen, zweierlei Recht. Das dürfte so viel böses Blut zur Folge haben, dass das auch nicht im Interesse der Vertriebenen liegen kann. Herr Kollege Dr. Kather hat vorhin gesagt, dass 300 000 bäuerliche Familien aus dem Osten geflüchtet sind. Infolge der Formulierung des § 35 ist die Zahl der Bewerber um einen Betrieb weitaus größer geworden. Ich meine, wir sind unehrlich, wenn wir den Leuten Hoffnungen machen, die doch nicht zu erfüllen sind. ? ?Thus even the third, fourth, and fifth son of an East German farmer […] can receive all the benefits. […] Which local farmer […] can earn such a dowry for three or four of his sons? If we restore the old version of the bill, we create two kinds of law. That would cause so much bad blood that it cannot possibly be in the interest of the expellees themselves. […] Dr. Kather said that 300,000 farming families fled from the East. Because of the wording of § 35, the number of applicants for a farm has become far greater. I believe we are being dishonest if we give people hopes that cannot in fact be fulfilled.?
Table 19: Examples for postwar period discussed in Section˜4.2 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(18) Dominant DSL subtype score: compassionate solidarity towards refugees